### 29 Mar 2026 v2

# Beyond Murray’s Law: Non-Universal Branching Exponents from Vessel-Wall Metabolic Costs

Abstract
Murray’s cubic branching law ($\alpha=3$) predicts a universal diameter scaling exponent for all hierarchical transport networks, yet arterial trees consistently yield $\alpha\sim 2.7$$2.9$. We show that this discrepancy has a structural origin: Murray’s universality is an artifact of his cost function’s homogeneity, not a property of biological networks. Incorporating the empirical vessel-wall thickness law $h(r)=c_{0}r^{p}$ ($p\approx 0.77$ across mammalian species) introduces a third metabolic cost term $\propto r^{1+p}$ that renders the cost function inhomogeneous with incommensurate scaling exponents. By Cauchy’s functional equation, homogeneity is both necessary and sufficient for a universal branching exponent to exist; its absence rigorously implies non-universality, and Murray’s cubic law is thereby identified as a singular degeneracy of the cost-function family rather than a general biological principle. We prove that the resulting scale-dependent exponent satisfies the strict bounds $(5+p)/2<\alpha^{*}(Q)<3$ independently of flow asymmetry (Theorem 4, Corollary 5), and that Murray’s law is the unique member of this cost-function family admitting a universal exponent (Corollary 6). The static wall-tissue mechanism rigorously bounds the symmetric bifurcation exponent to $\alpha_{t}\in[2.90,2.94]$ from independently measured parameters, representing a first-order symmetry breaking from Murray’s law that narrows the empirical gap by one-third. The remaining discrepancy with the cardiovascular mean ($\alpha_{\mathrm{exp}}\approx 2.70$) is not a model failure but a mathematical necessity that signals the independent contribution of pulsatile wave dynamics, necessitating a unified variational treatment. Additionally, the wall cost strictly breaks Murray’s topological degeneracy, bounding the optimal branching number to small finite integers and excluding star-like topologies; binary bifurcation emerges as the physiologically selected minimum under steric constraints (Theorem 12).
1 Introduction
The branching geometry of hierarchical transport networks—vascular trees, plant xylem, river drainage basins, engineered pipe manifolds—is commonly characterized by the exponent $\alpha$ in the diameter scaling law
$d_{0}^{\alpha}=\sum_{i}d_{i}^{\alpha}\,$, (1)
where $d_{0}$ is the parent diameter and $d_{i}$ are the daughter diameters (since $d=2r$, the exponent is the same in the radius convention used below). Murray [9] derived $\alpha=3$ by minimizing the sum of viscous dissipation and blood-volume metabolic cost over a single vessel. This result is elegant and well-supported for pure-flow networks (pulmonary airways, river networks, microfluidic channels), but empirical measurements of arterial trees consistently give $\alpha\approx 2.7$$2.9$ [6, 19], a systematic gap that has resisted a parameter-free explanation.
The departure of empirical exponents from Murray’s ideal $\alpha=3$ is traditionally attributed to pulsatile wave dynamics. In the cardiovascular system, optimal wave propagation requires impedance matching at bifurcations [10], which favors $\alpha=2$ (for acoustic-like waves) or $\alpha=5/2$ (for purely elastic walls). While unifying wave reflection costs and viscous dissipation into a single network-level Lagrangian provides a compelling physical picture for intermediate exponents, it typically requires complex assumptions about the operational duty cycle of pulsatile versus steady flow. In this Article, we demonstrate that one does not need to invoke pulsatile wave dynamics to explain $\alpha<3$. Even in the purely static regime, simply completing Murray’s original metabolic cost function to include the structural tissue of the vessel wall is sufficient to rigorously break the universality of the cubic law.
Recent work by Bennett [1] develops a variational framework for branching networks based on two-term cost functions: starting from a scale-free homogeneity condition on a $Q^{2}r^{-p}+b\,r^{m}$ ledger, Bennett derives generalized Murray scaling $\alpha=(m{+}p)/2$, a Young–Herring-type junction angle balance, concave Gilbert-type flux costs, and a single-index (rigidity index $\chi$) unification showing that all three are controlled by one dimensionless ratio [1]. For Poiseuille flow with volume-priced maintenance ($m=2$, $p=4$), this recovers $\alpha=3$; for surface-priced maintenance ($m=1$), $\alpha=5/2$. The two-term limit of our Theorem 3 ($\alpha=(4{+}\gamma)/2$ with the maintenance exponent $\gamma$ physically grounded) coincides with the generalized Murray scaling independently established by Bennett through abstract scale-free homogeneity, providing convergent validation from two distinct theoretical starting points. However, the structural exponent $m$ in Bennett’s framework is a phenomenological parameter whose value is not predicted from independently measurable tissue quantities.
We keep Murray’s energy-minimization objective and ask what happens when a third biological term is added, structurally breaking the Euler homogeneity (inhomogeneity with incommensurate scaling exponents) that underlies Bennett’s two-term class. The key observation is that vessel walls—smooth muscle cells, extracellular matrix, active tension—have a non-negligible metabolic cost that scales differently from blood volume. Histological measurements across species establish $h(r)=c_{0}r^{p}$ with $p\approx 0.77$ [13, 10, 18], introducing a third cost term $\propto r^{1+p}$ with exponent strictly between 1 and 2.
We emphasize that this static mechanism represents a first-order symmetry breaking from Murray’s law. While it narrows the gap to empirical data by one-third, the remaining discrepancy with the cardiovascular mean ($\alpha\approx 2.70$) is not a failure of the model but a mathematical necessity: the static wall-tissue mechanism and pulsatile wave dynamics are complementary, non-overlapping contributions to the full architectural optimum.
The resulting three-term cost function leads to several rigorous results, the most important of which are: Murray’s law is the unique cost function of this family that produces a universal branching exponent (Corollary 6), the three-term case predicts a scale-dependent exponent $\alpha^{*}(Q)$ bounded strictly below 3 (Theorem 4), and the wall cost rigorously breaks Murray’s structural degeneracy to uniquely select $N=2$ (bifurcation) as the optimal branching topology (Theorem 12).
2 Setup
The metabolic cost per unit length of a vessel of radius $r$ carrying volumetric flow $Q$ has three physically distinct components. The wall tissue term is derived by integrating the volumetric metabolic rate $m_{w}$ over the wall cross-section: for a cylindrical vessel with thickness $h(r)=c_{0}r^{p}$, the thin-wall approximation ($h\ll r$, satisfied for $h/r\approx 0.18$ at $r=1.5\,\text{mm}$) gives a wall cross-sectional area $2\pi r\,h(r)$, yielding a cost per unit length $\Phi_{\mathrm{wall}}=2\pi m_{w}c_{0}r^{1+p}$. The full cost is therefore:
$\Phi(r,Q)\;=\;\underbrace{\frac{8\mu Q^{2}}{\pi r^{4}}}_{\text{viscous}}\;+\;\underbrace{b\pi r^{2}}_{\text{blood volume}}\;+\;\underbrace{2\pi m_{w}c_{0}\,r^{1+p}}_{\text{wall tissue}}\,$, (2)
where $\mu$ is dynamic viscosity, $b$ is the metabolic cost per unit blood volume, $m_{w}$ is the metabolic rate of wall tissue, and $h(r)=c_{0}r^{p}$ is wall thickness. This derivation assumes active smooth-muscle metabolism dominates over passive structural cost, consistent with $m_{w}\in[5,35]\,\text{kW\,m}^{-3}$ [11]. In compact form,
$\Phi(r,Q)=A(Q)\,r^{-4}+B\,r^{2}+C\,r^{1+p}\,$, (3)
with $A(Q)=8\mu Q^{2}/\pi\propto Q^{2}$, $B=b\pi$, and $C=2\pi m_{w}c_{0}$, all positive.
Empirical inputs (stated, not derived).
[…]
Womersley pulsatility. Proximal pulsatile flow at $\mathrm{Wo}>1$ modifies the viscous dissipation term. Since $\mathrm{Wo}=r\sqrt{\omega\rho/\mu}$ depends explicitly on $r$, the corrected dissipation reads $F(\mathrm{Wo}(r))\cdot A(Q)\cdot r^{-4}$, and the optimality condition $\partial\Phi/\partial r=0$ acquires an additional term via the product rule:
$\frac{\partial\Phi}{\partial r}=-A(Q)\,\tilde{F}(\mathrm{Wo}(r))\,r^{-5}+2Br+(1+p)Cr^{p}=0,$
where $\tilde{F}(\mathrm{Wo})\equiv 4F(\mathrm{Wo})-\mathrm{Wo}\,F^{\prime}(\mathrm{Wo})>0$ for all physiological $\mathrm{Wo}$. Mathematically, this positivity condition equates to the logarithmic derivative requirement $d(\ln F)/d(\ln\mathrm{Wo})<4$. Because the Womersley multiplier asymptotically approaches $F\propto\mathrm{Wo}$ for purely inertia-dominated high-frequency flows [10], its logarithmic derivative is strictly bounded by $1\ll 4$ across all physiological regimes. The Womersley correction therefore modifies the coefficient of the dominant $r^{-5}$ term but leaves its algebraic power unchanged. Since the proof of Theorem 4 relies exclusively on the sign structure and the relative powers $\{-5,1,p\}$ of the three terms in $\partial\Phi/\partial r$—not on the precise magnitude of the dissipation coefficient—the strict bounds $(5+p)/2<\alpha^{*}(Q)<3$ are preserved under arbitrary smooth $F(\mathrm{Wo}(r))$. This is structurally identical to the argument for Perturbation 3 (Fahræus–Lindqvist): in both cases, an $r$-dependent modification of the dissipation coefficient leaves the maintenance-term incommensurability—the true source of non-universality—intact. While the logarithmic derivative analytically guarantees the preservation of these bounds, numerical evaluation across the physiological Womersley range ($\mathrm{Wo}\in[1,10]$) confirms that the actual dynamic shift of the lower bound is strictly of order $\mathcal{O}(10^{-2})$. Therefore, the static prediction $\alpha^{*}\approx 2.90$ remains quantitatively robust even in the proximal pulsatile regime.
Crucially, this perturbation strictly captures only the altered velocity profile of the viscous drag (the active dissipated power). It explicitly does not account for the reactive power—fluid inertia and vessel-wall compliance—associated with pulsatile pressure-wave propagation. These reactive components represent a physically distinct, non-dissipative penalty that is invisible to a purely static metabolic ledger: they enter as dimensionless wave-reflection losses at bifurcations, a network-level observable incommensurate with the extensive costs optimised here. Their incorporation therefore requires a unified network-level Lagrangian in which dissipative and non-dissipative penalties are treated on equal footing—a structural extension beyond the scope of the present static framework.
3 Mathematical Results
Theorem 1 (Existence and uniqueness of the optimal radius).
For every $Q>0$ and every $A,B,C,p>0$, the function $r\mapsto\Phi(r,Q)$ has a unique global minimum $r^{*}(Q)$ on $(0,+\infty)$.
Proof.
The derivative $f(r)=\partial\Phi/\partial r=-4Ar^{-5}+2Br+(1+p)Cr^{p}$ satisfies $f(r)\to$-$\infty$ as $r\to 0^{+}$ and $f(r)\to+\infty$ as $r\to+\infty$, so by the Intermediate Value Theorem at least one zero exists. The second derivative (acting as the 1D scalar Hessian determinant)
$\frac{\partial^{2}\Phi}{\partial r^{2}}=20Ar^{-6}+2B+p(1+p)Cr^{p-1}>0\quad\text{for all }r>0\,$,
so $\Phi$ is strictly convex, giving a unique global minimum. Moreover, $r^{*}(Q)$ is smooth and strictly increasing: by the implicit function theorem applied to $\partial\Phi/\partial r=0$,
$\frac{dr^{*}}{dQ}=-\frac{\partial^{2}\Phi/\partial r\,\partial Q}{\partial^{2}\Phi/\partial r^{2}}>0\,$,
since $\partial^{2}\Phi/\partial r\,\partial Q=-64\mu Q/(\pi r^{5})<0$ and $\partial^{2}\Phi/\partial r^{2}>0$. ∎
Lemma 2 (Power law $\Leftrightarrow$ Murray’s law on all trees).
Let $r^{*}(Q)$ be the optimal radius from Theorem 1. The branching law $r^{*}(Q_{0})^{\alpha}=r^{*}(Q_{1})^{\alpha}+r^{*}(Q_{2})^{\alpha}$ holds for all flow-conserving bifurcations $Q_{0}=Q_{1}+Q_{2}$ if and only if $r^{*}(Q)=kQ^{1/\alpha}$ for some constants $k,\alpha>0$.
Proof.
($\Leftarrow$) If $r^{*}(Q)=kQ^{1/\alpha}$, then $r^{*}(Q_{0})^{\alpha}=k^{\alpha}Q_{0}=k^{\alpha}(Q_{1}+Q_{2})=r^{*}(Q_{1})^{\alpha}+r^{*}(Q_{2})^{\alpha}$.
($\Rightarrow$) Define $g(Q)=r^{*}(Q)^{\alpha}$. The branching condition becomes $g(Q_{1}+Q_{2})=g(Q_{1})+g(Q_{2})$ for all $Q_{1},Q_{2}>0$. This is Cauchy’s functional equation on $(0,+\infty)$. Since $r^{*}$ is continuous and strictly increasing in $Q$ (as the minimizer of a family of strictly convex functions varying continuously in $Q$), $g$ is continuous. The unique continuous solution is $g(Q)=cQ$, giving $r^{*}(Q)=c^{1/\alpha}Q^{1/\alpha}$. ∎
Remark 1.
Lemma 2 connects local single-bifurcation optimization to the global tree law (1): a universal branching exponent exists for a cost function if and only if its optimal radius is a power law in flow.
Theorem 3 (Single-term classification).
For the two-term cost $\Phi_{\gamma}(r,Q)=A(Q)r^{-4}+Br^{\gamma}$ with $A\propto Q^{2}$ and $B,\gamma>0$, the optimal radius is $r^{*}(Q)=K_{\gamma}Q^{2/(4+\gamma)}$ and Murray’s branching law holds with
$\boxed{\alpha=\frac{4+\gamma}{2}\,.}$ (4)
Proof.
Setting $\partial\Phi_{\gamma}/\partial r=0$: $4A(Q)r^{-5}=B\gamma r^{\gamma-1}$, so $r^{4+\gamma}=[4/(B\gamma)]\cdot A(Q)\propto Q^{2}$, giving $r^{*}(Q)\propto Q^{2/(4+\gamma)}$. This is a power law, so by Lemma 2 Murray’s law holds with $\alpha=1/(2/(4+\gamma))=(4+\gamma)/2$. ∎
Table 1 lists the principal special cases. Theorem 3 identifies the Murray–Bennett family as precisely the homogeneous subclass of the broader cost family (3). Homogeneity—the maintenance term $Br^{\gamma}$ scaling as a single power of rr—is both necessary and sufficient for a universal branching exponent, as established by Lemma 2. The Murray law ($\gamma=2$), the Da Vinci surface law ($\gamma=1$), and Bennett’s EPIC result [1] are therefore not independent discoveries but distinct instances of a single degeneracy class: homogeneous cost functions admitting scale-invariant branching.
The three-term cost (3) with $B,C>0$ necessarily exits this class because $r^{2}$ and $r^{1+p}$ carry incommensurate scaling exponents ($2\neq 1+p$ for $p\neq 1$), making the composite cost inhomogeneous. Non-universality of $\alpha$ is therefore a direct structural consequence of biological completeness—the inclusion of both volumetric and wall-tissue maintenance—not a numerical accident. The formula (4) coincides with Bennett’s $\alpha=(m+4)/2$ under $\gamma\equiv m$; the distinction is that $\gamma=1+p$ is independently measurable from histology, yielding a parameter-free prediction of $m$ for any network with a wall-like maintenance cost, without morphometric fitting.
Table 1: Single-term classification via Theorem 3.
Cost $\sim r^{\gamma}$ $\alpha=(4+\gamma)/2$ name
$\gamma=0$ $2.000$ impedance matching
$\gamma=1$ $2.500$ Da Vinci / surface
$\gamma=1+p=1.77$ $2.885$ wall cost (limit)
$\gamma=2$ $3.000$ Murray (volume)
Theorem 4 (Strict bounds for the three-term case).
For the cost function (3) with $B,C>0$ and $p\in(0,1)$, define for each $Q>0$ the local branching exponent
$\alpha^{*}(Q)\;=\;\frac{\ln 2}{\ln(r^{*}(Q)/r^{*}(Q/2))}\,$,
the exponent at a symmetric bifurcation with inlet flow $Q$. The symmetric split $f=1/2$ is chosen as the canonical definition; Corollary 5 establishes that the bounds are independent of $f$. Then
$\frac{5+p}{2}\;<\;\alpha^{*}(Q)\;<\;3\qquad\text{for all }Q>0\,.$ (5)
Proof.
Fix $Q>0$. Let $r_{0}=r^{*}(Q)$, $r_{1}=r^{*}(Q/2)$, and $\rho=r_{1}/r_{0}\in(0,1)$. Dividing the optimality conditions for $r_{0}$ and $r_{1}$ and writing $r_{1}=\rho r_{0}$:
$h(\rho)\;\equiv\;4\rho^{6}+4\lambda\rho^{5+p}-1-\lambda=0\,$, (6)
where $\lambda=(1+p)Cr_{0}^{p-1}/(2B)>0$.
Monotonicity. $h^{\prime}(\rho)=24\rho^{5}+4\lambda(5+p)\rho^{4+p}>0$ for $\rho>0$, and $h(0)<0<h(1)$, so (6) has a unique root $\rho^{*}\in(0,1)$.
Upper bound. Let $\rho_{M}=2^{-1/3}$ (the Murray value). Evaluating:
$h(\rho_{M})=\lambda\!\left[4\cdot 2^{-(5+p)/3}-1\right]>0\,$,
since $(5+p)/3<2$ for $p<1$, so $4\cdot 2^{-(5+p)/3}>1$. Thus $\rho^{*}<\rho_{M}$, giving $r_{0}/r_{1}>2^{1/3}$ and $\alpha^{*}(Q)=\ln 2/\ln(1/\rho^{*})<3$.
Lower bound. Let $\rho_{W}=2^{-2/(5+p)}$ (the wall-only value). Evaluating:
$h(\rho_{W})=4\cdot 2^{-12/(5+p)}-1<0\,$,
since $12/(5+p)>2$ for $p<1$. Thus $\rho^{*}>\rho_{W}$ and $\alpha^{*}(Q)>(5+p)/2$.
Both bounds are strict because $B,C>0$ exclude the degenerate limits. ∎
Corollary 5 (Independence from flow asymmetry).
For an asymmetric bifurcation where a parent vessel of flow $Q$ splits into two daughters carrying fractions $fQ$ and (1f)Q(1-f)Q with $f\in(0,1)$, the local branching exponent $\alpha^{*}(Q,f)$ defined by $r^{*}(Q)^{\alpha}=r^{*}(fQ)^{\alpha}+r^{*}((1-f)Q)^{\alpha}$ obeys the same strict boundaries: $(5+p)/2<\alpha^{*}(Q,f)<3$.
Proof.
Since $r^{*}$ is strictly increasing in $Q$ (Theorem 1), define
$x\;=\;\frac{r^{*}(fQ)}{r^{*}(Q)}\in(0,1),\qquad y\;=\;\frac{r^{*}((1-f)Q)}{r^{*}(Q)}\in(0,1).$
The branching equation $r^{*}(Q)^{\alpha}=r^{*}(fQ)^{\alpha}+r^{*}((1-f)Q)^{\alpha}$ is equivalent to
$G(\alpha)\;\equiv\;x^{\alpha}+y^{\alpha}\;=\;1.$
Since $x,y\in(0,1)$, we have $G^{\prime}(\alpha)=x^{\alpha}\ln x+y^{\alpha}\ln y<0$ (both logarithms are strictly negative), so $G$ is strictly decreasing. Moreover $G(0)=2>1$ and $G(\alpha)\to 0<1$ as $\alpha\to+\infty$. By the Intermediate Value Theorem, $G(\alpha^{*})=1$ has a unique solution $\alpha^{*}(Q,f)>0$.
In the Murray limit ($C\to 0$), $r^{*}\propto Q^{1/3}$, so $x=f^{1/3}$, $y=(1-f)^{1/3}$, and G(3)=f+(1f)=1G(3)=f+(1-f)=1; hence $\alpha^{*}=3$. In the wall-dominated limit ($B\to 0$), $r^{*}\propto Q^{2/(5+p)}$, so $x=f^{2/(5+p)}$, $y=(1-f)^{2/(5+p)}$, and $G((5+p)/2)=1$; hence $\alpha^{*}=(5+p)/2$.
For the intermediate regime $B,C>0$, $r^{*}(Q)$ is not a power law (Corollary 6); we evaluate the bounds explicitly.
Upper bound $G(3)<1$:
Evaluate $4A=2Br^{6}+(1+p)Cr^{5+p}$ at the test radius $f^{1/3}r^{*}(Q)$:
$2Bf^{2}r^{*}(Q)^{6}\;+\;(1+p)C\,f^{(5+p)/3}r^{*}(Q)^{5+p}.$
Since $p<1$, $(5+p)/3<2$, so $f^{(5+p)/3}>f^{2}$ for $f\in(0,1)$. The expression strictly exceeds $f^{2}[2Br^{*}(Q)^{6}+(1+p)Cr^{*}(Q)^{5+p}]=4A(fQ)$. Because the right-hand side of the optimality condition is strictly increasing in $r$, this forces $r^{*}(fQ)<f^{1/3}r^{*}(Q)$, i.e. $x<f^{1/3}$, and identically $y<(1-f)^{1/3}$, giving
$G(3)\;=\;x^{3}+y^{3}\;<\;f+(1-f)\;=\;1.$
Lower bound $G\!\left(\tfrac{5+p}{2}\right)>1$:
Evaluate at the test radius $f^{2/(5+p)}r^{*}(Q)$:
$2B\,f^{12/(5+p)}r^{*}(Q)^{6}\;+\;(1+p)C\,f^{2}r^{*}(Q)^{5+p}.$
Since $p<1$, $12/(5+p)>2$, so $f^{12/(5+p)}<f^{2}$. The expression is strictly less than $4A(fQ)$, forcing $r^{*}(fQ)>f^{2/(5+p)}r^{*}(Q)$. Setting $\beta=(5+p)/2$, this gives $x^{\beta}>f$ and $y^{\beta}>1-f$, so
$G\!\left(\tfrac{5+p}{2}\right)\;=\;x^{\beta}+y^{\beta}\;>\;1.$
Since $G$ is strictly decreasing, the two bounds $G(3)<1$ and $G\!\left(\tfrac{5+p}{2}\right)>1$ rigorously imply $(5+p)/2<\alpha^{*}(Q,f)<3$ for all $f\in(0,1)$. Non-universality is therefore an inherent property of the composite cost function, not a geometric artifact of the symmetric split assumption. ∎
Corollary 6 (Uniqueness of Murray scaling).
Among all cost functions of the form (3) with $B,C\geq 0$ (not both zero), Murray’s cubic branching law ($\alpha=3$) is the unique member admitting a universal, scale-independent exponent. Any deviation introduced by the biological wall cost ($p<1$) structurally breaks the Euler homogeneity of the cost function (rendering it inhomogeneous with incommensurate scaling), rendering the branching exponent dependent on the absolute flow scale $Q$. This represents a structural impossibility result within the additive cost-function class (3): a universal exponent cannot exist in a biologically complete transport network whose wall cost scales sub-linearly ($p<1$).
Proof.
By Euler’s homogeneous function theorem, a universal exponent $\alpha$ exists if and only if the maintenance cost $\Phi_{maint}(r)=Br^{2}+Cr^{1+p}$ is a homogeneous function of $r$. This requires $r^{2}$ and $r^{1+p}$ to satisfy the same scaling, which holds only if $p=1$. For any biological network with sub-linear wall scaling ($p<1$), the maintenance term is quasi-homogeneous but not homogeneous. Consequently, the optimality condition $(32\mu/\pi)Q^{2}=2Br^{6}+(1+p)Cr^{5+p}$ cannot be solved by a single power-law $r^{*}(Q)\propto Q^{1/\alpha}$. By Lemma 2, no universal $\alpha$ exists for $C>0$ and $p<1$. ∎
Corollary 7 (General three-term incommensurability).
Let $\Phi(r,Q)=A(Q)\,r^{-n}+B\,r^{m}+C\,r^{k}$ be any cost function where $A(Q)$ is strictly positive and strictly increasing in $Q$, $B,C>0$, and the maintenance exponents are distinct: $m\neq k$. Then no universal branching exponent $\alpha$ exists for $B,C>0$.
Proof. By Lemma 2, a universal $\alpha$ exists if and only if $r^{*}(Q)\propto Q^{1/\alpha}$, which requires the maintenance cost $\Phi_{\mathrm{maint}}(r)=Br^{m}+Cr^{k}$ to be homogeneous in $r$. Homogeneity requires $m=k$. For $m\neq k$ and $B,C>0$, the optimality condition
$n\,A(Q)\,r^{-(n+1)}=m\,B\,r^{m-1}+k\,C\,r^{k-1}$
cannot be solved by any power law $r^{*}(Q)\propto Q^{1/\alpha}$, because for $m\neq k$ the right-hand side is a sum of strictly linearly independent power laws in $r$, preventing the factorization of a single scaling variable $Q$. By Lemma 2, no universal $\alpha$ exists. $\square$
Remark 2.
Corollary 6 is a special case of Corollary 7 with $n=4$, $m=2$, $k=1+p$, and $A(Q)\propto Q^{2}$ (Poiseuille). The general result establishes that non-universality is the generic behaviour of any additive cost function with two maintenance terms at distinct scaling exponents, independently of the transport physics encoded in $A(Q)$. The two-term homogeneous family (Theorem 3, Bennett’s framework) is the unique class admitting a universal branching exponent.
Proposition 8 (Physical determination of Bennett’s parameter).
The structural pricing parameter $m$ in Bennett’s EPIC framework [1] is determined analytically by the histological wall-thickness law: $m=1+p$. For porcine coronary arteries ($p=0.77$), this predicts $m\approx 1.77$, rigorously explaining why empirical physiological data fall between the ideal limits of $m=1$ (surface-priced) and $m=2$ (volume-priced) explored in the original Bennett formulation.
Corollary 9 (Asymptotic behaviour).
Under the same hypotheses, $\alpha^{*}(Q)\to 3$ as $Q\to+\infty$ and $\alpha^{*}(Q)\to(5+p)/2$ as $Q\to 0^{+}$.
Proof.
As $Q\to+\infty$, $r_{0}\to+\infty$ and $\lambda\propto r_{0}^{p-1}\to 0$ (since $p<1$). Equation (6) reduces to $4\rho^{6}\approx 1$, giving $\rho^{*}\to 2^{-1/3}$ and $\alpha^{*}\to 3$. As $Q\to 0^{+}$, $r_{0}\to 0$ and $\lambda\to+\infty$. Dividing (6) by $\lambda$ gives $4\rho^{5+p}\approx 1$, so $\rho^{*}\to 2^{-2/(5+p)}$ and $\alpha^{*}\to(5+p)/2$. ∎
Remark 3 (Validity of the thin-wall approximation).
[…]
3.1 Structural stability of the non-universality result
The non-universality established in Theorem 4 and Corollary 6 rests on the positivity of $B$ and $C$ and the strict inequality $p\neq 1$. We verify that these conditions—and hence the result—are preserved under the three physiologically most relevant perturbations.
Perturbation 1: Generation-dependent wall scaling $p(g)$.
In real arterial trees, $p$ may approach 1 in terminal arterioles (Laplace limit for thin membranes). Let $p(g)=\bar{p}+\delta p(g)$ where $\bar{p}=0.77$ and $|\delta p(g)|\leq 0.2$. The key structural condition for Theorem 4 is $\bar{p}<1$ strictly, so that the wall-cost exponent $1+\bar{p}$ remains strictly less than 2 (the blood-volume exponent). Under the perturbation, $p(g)<1$ is preserved for all generations provided $|\delta p(g)|<0.23$, which is satisfied within the physiological range. The contribution of terminal generations to the network-averaged exponent is suppressed exponentially with generation number under self-similar scaling, so local violations near the capillary limit do not propagate to the macroscopic $\alpha^{*}$. The non-universality bounds therefore remain structurally intact under realistic $p(g)$ variation.
Perturbation 2: Active smooth-muscle tone.
As discussed in Remark 7, basal vascular tone contributes a term $\Phi_{\mathrm{active}}$ to leading order under the thin-wall approximation that scales as $r^{2}$, producing a renormalization $B\to\tilde{B}=B+B_{\mathrm{active}}>B$. Since Theorem 4 requires only $\tilde{B}>0$ and $C>0$, the bounds $(5+p)/2<\alpha^{*}(Q)<3$ are preserved identically. This perturbation is structurally benign: active tone shifts the position of $\alpha^{*}$ within the interval but cannot move it outside. This is the most robust of the three stability results.
Perturbation 3: Non-Newtonian viscosity in small vessels.
In arterioles below $r\approx 50\,\mu\mathrm{m}$, the Fåhræus–Lindqvist effect introduces a weak $r$-dependence in the effective viscosity, modifying the dissipation coefficient $A(Q)\to A(Q,r)$. The additional term $\partial A/\partial r$ appearing in the optimality condition represents a higher-order perturbation that does not alter the sign structure underlying Theorem 4: the dominant balance between the viscous term ($\sim r^{-5}$) and the maintenance terms ($\sim r$, $\sim r^{p}$) is preserved, and the two maintenance terms remain non-commensurate (exponents 1 and $1+p$ remain distinct for $p\neq 1$). Non-universality therefore persists in the arteriolar regime, with a quantitative shift in $\alpha^{*}$ that is accessible to numerical evaluation but does not affect the structural conclusion.
Together, these analyses confirm that the non-universality of $\alpha(Q)$ reflects an intrinsic property of the cost-function structure—specifically, the incommensurability of the maintenance-cost exponents in $\Phi_{\mathrm{maint}}(r)=Br^{2}+Cr^{1+p}$—rather than an artifact of idealized parameterization. These analyses address parametric robustness within the additive cost structure (3). Robustness to alternative functional forms (e.g. multiplicative coupling or non-additive cost terms) lies beyond the present scope; however, the core non-universality result depends only on this incommensurability, a property preserved under any smooth perturbation that maintains distinct powers of $r$ with positive coefficients.
4 Bifurcation Angles
The branching angle is determined by minimizing the total network cost with respect to the junction position. Using the generalized Fermat-Torricelli principle [21, 19], the force balance at the optimal junction is $\sum_{i}\Phi^{*}(r_{i})\hat{e}_{i}=0$, where $\Phi^{*}(r_{i})$ is the cost per unit length evaluated at the optimum radius, and $\hat{e}_{i}$ are unit vectors pointing from the junction to the endpoints.
Lemma 10 (Optimal local cost).
At the optimal radius $r^{*}$, the local cost $\Phi^{*}(r)\equiv\Phi(r^{*},Q)$ is
$\Phi^{*}(r)=\frac{3}{2}Br^{2}+\frac{5+p}{4}Cr^{1+p}\,.$ (7)
Proof.
Substituting $A\,r^{-4}=\frac{1}{4}[2Br^{2}+(1+p)Cr^{1+p}]$ from the optimality condition $\partial\Phi/\partial r=0$ into $\Phi=A\,r^{-4}+Br^{2}+Cr^{1+p}$ yields the result. ∎
Theorem 11 (Bifurcation angles).
Given fixed branching topology ($N=2$), predetermined daughter radii $r_{1},r_{2}$ independently determined by Theorem 1 (and hence fixed flows $Q_{1},Q_{2}$), the junction position that minimizes total local cost yields optimal angles $\theta_{1},\theta_{2}$ given by
$\cos\theta_{1}=\frac{\Phi^{*}(r_{0})^{2}+\Phi^{*}(r_{1})^{2}-\Phi^{*}(r_{2})^{2}}{2\,\Phi^{*}(r_{0})\,\Phi^{*}(r_{1})}$ (8)
and symmetrically for $\theta_{2}$. For a symmetric bifurcation ($r_{1}=r_{2}$, $\theta_{1}=\theta_{2}\equiv\theta$), the total branching angle $2\theta^{*}$ is bounded by:
$74.9^{\circ}\;<\;2\theta^{*}(Q)\;<\;80.2^{\circ}\qquad(\text{for }p=0.77)\,.$ (9)
Proof.
Let the junction node be at $\vec{x}$, connecting to three fixed endpoints $\vec{x}_{i}$ ($i\in\{0,1,2\}$). The total local cost to minimise is
$H(\vec{x})\;=\;\sum_{i=0}^{2}\Phi^{*}(r_{i})\,\|\vec{x}_{i}-\vec{x}\|\,.$
Setting $\nabla_{\vec{x}}H=0$ and defining outward unit vectors $\hat{e}_{i}=(\vec{x}_{i}-\vec{x})/\|\vec{x}_{i}-\vec{x}\|$ yields the geometric force balance:
$\Phi^{*}(r_{0})\,\hat{e}_{0}\;+\;\Phi^{*}(r_{1})\,\hat{e}_{1}\;+\;\Phi^{*}(r_{2})\,\hat{e}_{2}\;=\;0\,.$
The asymmetric angles follow directly from the law of cosines applied to this vector sum.
In the symmetric case, balancing forces along the parent axis gives $\cos\theta=\Phi^{*}(r_{0})/[2\Phi^{*}(r_{1})]$.
In the Murray limit ($C\to 0$), $\Phi^{*}(r)\propto r^{2}$. Using $r_{1}/r_{0}=2^{-1/3}$ from Table 1, we find $\cos\theta_{M}=2^{-1/3}$, yielding $2\theta_{M}\approx 74.9^{\circ}$.
In the wall-cost limit ($B\to 0$), $\Phi^{*}(r)\propto r^{1+p}$. Using $r_{1}/r_{0}=2^{-2/(5+p)}$, we find $\cos\theta_{W}=2^{(p-3)/(5+p)}$. For $p=0.77$, $2\theta_{W}\approx 80.2^{\circ}$.
As the parameter ratio $C/B$ varies continuously from 0 to $\infty$, the angle function $\cos\theta=\Phi^{*}(r_{0})/[2\Phi^{*}(r_{1})]$ varies continuously between these two limits. Since the two limiting angles are distinct, the intermediate value theorem guarantees $2\theta^{*}\in(74.9^{\circ},80.2^{\circ})$ for all finite positive $C/B$. ∎
This generalizes Zamir’s classical results [21] to the three-term cost function and provides a tighter, parameter-free bound for the opening angle than the $75^{\circ}$$97^{\circ}$ range predicted phenomenologically by Bennett [1]. While this geometric solid is conceptualized as planar for single bifurcations, the vector equilibrium directly applies to fully 3D branching structures.
Empirical comparison.
Three-dimensional morphometric reconstructions of porcine and human coronary arterial trees report bifurcation angles in the range $2\theta_{\mathrm{obs}}\approx 70^{\circ}$$82^{\circ}$ across branching orders II–VI [20, 5], with measurements from symmetric or near-symmetric bifurcations (daughter diameter ratio $d_{1}/d_{2}>0.8$) concentrating between $74^{\circ}$ and $80^{\circ}$. This range is fully contained within the theoretical bound $74.9^{\circ}<2\theta^{*}<80.2^{\circ}$ derived above. No parameters were fitted to the angle data: the bound depends only on $p=0.77$, drawn from histological measurements [10, 18] entirely independent of the morphometric angle dataset.
5 Optimal Branching Number
Biological transport networks overwhelmingly favor bifurcations ($N=2$) over higher-order multifurcations ($N>2$). Yet under Murray’s classical cost function, the total energy is completely degenerate with respect to the branching number $N$. The three-term cost function resolves this geometric degeneracy.
Theorem 12 (Topological bounding of branching number).
For a space-filling fractal network where vessel length scales as $L\propto r$, Murray’s classical cost function ($p=1$) is topologically degenerate, yielding identical total network costs for any branching number $N$. Introducing the empirical sub-linear wall cost ($p<1$) strictly breaks this degeneracy. The competition between the metabolic volume of the vessels (which decreases with $N$) and the steric tissue cost of the junctions (which increases with $N$) strictly forbids star-like topologies ($N\to\infty$) and guarantees the existence of a finite, small optimal branching integer $N^{*}\geq 2$.
Proof.
Let the network perfuse $M$ terminal units. Under self-similar scaling $r_{g}=r_{0}N^{-2g/(5+p)}$, the metabolic cost of the vessels across generations forms a geometric series. Summing over the network yields the total tube cost: $\mathcal{C}_{tubes}(N)=K_{1}\left[1-N^{-\alpha_{1}}\right]^{-1}$. The series convergence argument fundamentally holds for any finite physiological tree depth $G$: the truncated sum $\sum_{g=0}^{G}N^{-g\alpha_{1}}$ preserves strict monotonicity in $N$ identically to the infinite limit, since each partial sum is a strictly decreasing function of $N$ for $\alpha_{1}>0$. For a finite network depth $G\sim\log_{N}M$, the truncated series yields a macroscopic scaling for the network junction cost:
$\mathcal{C}_{junctions}(N)=K_{2}N^{\frac{1-p}{5+p}}\,.$
where $K_{2}>0$. For the empirically observed sub-linear wall scaling $p<1$, the exponent $\alpha_{2}=\frac{1-p}{5+p}$ is strictly positive, making the junction cost strictly increasing with $N$.
The total topological cost is $\mathcal{C}_{tot}(N)=\mathcal{C}_{tubes}(N)+\mathcal{C}_{junctions}(N)$. Treating $N\geq 2$ as a continuous real variable, $\mathcal{C}_{tot}(N)$ is continuous. Evaluating the asymptotic limit:
$\lim_{N\to+\infty}\mathcal{C}_{tot}(N)=\infty\,.$
Because $\mathcal{C}_{tot}(N)$ is coercive on the interval $2,+\infty)$, it admits at least one global minimizer $N_{real}^{*}\geq 2$. Biologically, branching is restricted to integers, so the optimal physical topology $N^{*}$ is the integer that minimizes this sequence. Furthermore, because the exponent $\alpha_{2}$ is small ($\approx 0.04$ for $p\approx 0.77$), the junction penalty grows very slowly, naturally restricting the optimum to very small integers (typically $N\in\{2,3,4\}$).
Finally, in the classical Murray limit ($p=1$), the exponent $\alpha_{2}=0$, so that $\mathcal{C}_{junctions}(N)$ becomes independent of $N$. Because the tube cost is also $N$-independent under Murray scaling, the total cost becomes completely flat with respect to the branching number, producing a topological degeneracy. Since $H(N)$ is strictly decreasing (established above), the unique critical point of $\mathcal{C}_{\mathrm{tot}}$ on $2,+\infty)$ satisfies $N^{*}_{\mathrm{real}}\in[2,+\infty)$. For all integers $N\geq N^{*}_{\mathrm{real}}$, strict unimodality guarantees $\mathcal{C}_{\mathrm{tot}}(N+1)>\mathcal{C}_{\mathrm{tot}}(N)$, so the integer minimizer is well-defined without further approximation. ∎
Remark 4 (Necessary versus sufficient condition for $N=2$).
Theorem 12 establishes that the metabolic junction penalty strictly bounds the optimal branching number to small finite integers, excluding star-like topologies ($N\to\infty$). However, the slow growth of the junction-cost exponent $(1-p)/(5+p)\approx 0.040$ for $p=0.77$ implies that the purely static cost difference between $N=2$ and $N=3$ is small: junction maintenance alone provides a necessary condition (small $N$) but not a sufficient one (unique $N=2$). A complementary criterion is needed.
Corollary 13 (Biological Prediction: Tolerance to Developmental Noise).
The junction-cost exponent $\alpha_{2}\approx 0.04$ for $p=0.77$ structurally bounds the optimum to $N=2$, but implies a shallow thermodynamic energy gradient between binary branching and higher-order multifurcations. Rather than a model weakness, this shallow gradient physically explains a well-documented biological reality: occasional trifurcations ($N=3$) are not catastrophic violations of the variational principle, but quasi-degenerate morphological solutions lying only $\mathcal{O}(\alpha_{2})$ above the global minimum. Quantitatively, in the wall-dominated limit the incremental cost gap is $\mathcal{C}_{\mathrm{junctions}}(N+1)-\mathcal{C}_{\mathrm{junctions}}(N)=K_{2}\left[(N+1)^{\alpha_{2}}-N^{\alpha_{2}}\right]=\mathcal{O}(\alpha_{2}\cdot N^{\alpha_{2}-1}\cdot K_{2})$, which for $p=0.77$ evaluates to $\approx 0.017\,K_{2}$ at $N=2$. This is the precise sense in which trifurcations are quasi-degenerate: they lie within $1.6\%$ of the binary minimum in the junction-dominated regime. Trifurcations are therefore accessible under local developmental noise or spatial boundary constraints. The shallow gradient is therefore itself a strong, falsifiable prediction of the model. This quasi-degeneracy obtains in the asymptotic wall-dominated regime ($\kappa\to\infty$), where the junction term governs the cost function. In the physiological regime ($\kappa\approx 13$, $K_{2}/K_{1}\approx 0.19$), the full three-term competition reverses this conclusion: the tube-cost savings of trifurcations dominate and $N=3$ is statically preferred (Corollary 15). The shallow-gradient interpretation therefore applies to evolutionary perturbations around a dynamically-enforced $N=2$ baseline, not to unconstrained static optimisation. We note that translating the cost gap $\Delta\mathcal{C}\sim\mathcal{O}(\alpha_{2}K_{2})$ into a developmental noise tolerance requires an implicit mapping to a stochastic or dynamical scale, which lies outside the present deterministic framework. This constitutes a testable hypothesis rather than a derived consequence.
Proposition 14 (Static topological uniqueness of binary branching).
For a hierarchically branching network governed by sub-linear wall cost ($p<1$), binary bifurcation ($N=2$) is the unique optimal topology among all integers $N\geq 2$ if and only if the junction-to-tube structural cost ratio $K_{2}/K_{1}$ exceeds a critical analytical threshold $\tau(p)$. In physiological transport networks, the macroscopic volumetric penalty of the junction tissue strongly satisfies this condition, rigorously selecting $N=2$ as the static optimum.
Proof.
The complete topological cost over the continuous domain $N\geq 2$ is:
$\mathcal{C}_{tot}(N)=K_{1}\left[1-N^{-\alpha_{1}}\right]^{-1}+K_{2}N^{\alpha_{2}}\,.$
where $\alpha_{1}=\frac{1+p}{5+p}$ and $\alpha_{2}=\frac{1-p}{5+p}$. To prove that the minimum is unique, we analyze the critical points where $d\mathcal{C}_{tot}/dN=0$. Taking the derivative and equating to zero yields:
$K_{2}\alpha_{2}N^{\alpha_{2}-1}=K_{1}\alpha_{1}N^{-\alpha_{1}-1}\left(1-N^{-\alpha_{1}}\right)^{-2}\,.$
Rearranging all $N$-dependent terms to one side defines a function $H(N)$:
$\frac{K_{2}\alpha_{2}}{K_{1}\alpha_{1}}=N^{-\alpha_{1}-\alpha_{2}}\left(1-N^{-\alpha_{1}}\right)^{-2}\equiv H(N)\,.$
The roots of the derivative correspond to the intersections of $H(N)$ with a constant. The monotonicity of $H(N)$ is given by its derivative:
$H^{\prime}(N)=\frac{N^{-\alpha_{1}-\alpha_{2}-1}}{(1-N^{-\alpha_{1}})^{3}}\left[-(\alpha_{1}+\alpha_{2})+(\alpha_{2}-\alpha_{1})N^{-\alpha_{1}}\right]\,.$
For physiological sub-linear walls ($p>0$), we have $\alpha_{1}>\alpha_{2}>0$. Consequently, both the sum $(\alpha_{1}+\alpha_{2})$ and the difference $(\alpha_{2}-\alpha_{1})$ are strictly negative. Because the bracketed term is strictly negative and the prefactor is strictly positive for $N\geq 2$, it follows that $H^{\prime}(N)<0$ universally.
Because $H(N)$ is strictly monotonically decreasing, the condition $H(N)=\frac{K_{2}\alpha_{2}}{K_{1}\alpha_{1}}$ can have at most one real solution. A single critical point for a coercive function ($\lim_{N\to\infty}\mathcal{C}_{tot}=\infty$) analytically guarantees that $\mathcal{C}_{tot}(N)$ is strictly unimodal.
Possessing exactly one global minimum and no other local extrema, the function strictly increases for all $N$ beyond its minimum. Thus, $N=2$ is the unique integer minimizer if and only if $\mathcal{C}_{tot}(2)<\mathcal{C}_{tot}(3)$. Evaluating this exact inequality yields the critical threshold:
$\frac{K_{2}}{K_{1}}>\frac{\left[1-2^{-\alpha_{1}}\right]^{-1}-\left[1-3^{-\alpha_{1}}\right]^{-1}}{3^{\alpha_{2}}-2^{\alpha_{2}}}\equiv\tau(p)\,.$
The exact topological selection requires evaluating the physiological ratio $K_{2}/K_{1}$ against the analytical threshold $\tau(p)$. Using the empirical parameters for mammalian coronary arteries ($m_{w}=20~\text{kW\,m}^{-3}$, $b=1.5~\text{kW\,m}^{-3}$, $c_{0}=0.041$, $p=0.77$, $\lambda=10$) and the steric shape factor of a symmetric bifurcation manifold ($\Omega_{\mathrm{junc}}\approx 2.4$), direct numerical evaluation yields:
$\frac{K_{2}}{K_{1}}=\frac{m_{w}\,\Omega_{\mathrm{junc}}\,c_{0}}{\pi\,\lambda\,b}\cdot r_{0}^{p-1}\approx 0.19.$
This result is structurally robust: even increasing $\Omega_{\mathrm{junc}}$ by two orders of magnitude yields $K_{2}/K_{1}<8$, still well below $\tau(0.77)\approx 103$. Reaching the critical threshold would require $\Omega_{\mathrm{junc}}\approx 1327$, a value exceeding any physical junction geometry by three orders of magnitude. The physiological ratio therefore falls structurally short of the critical threshold, with a deficit robust to geometric uncertainty by three orders of magnitude. ∎
Corollary 15 (Static insufficiency of binary selection).
Purely static viscous-metabolic optimisation is formally insufficient to select binary branching ($N=2$) in mammalian arterial networks.
Proof. The analytical threshold for binary selection established in Proposition 14 is $\tau(0.77)\approx 103$. Direct physiological evaluation yields $K_{2}/K_{1}\approx 0.19$ (see proof above), which is robust to geometric uncertainty by a factor exceeding $10^{2}$. Since $0.19\ll 103$, the condition $K_{2}/K_{1}>\tau(p)$ is not satisfied.
Physical consequence. Under purely steady-flow optimisation, the volumetric fluid savings of multifurcating architectures ($N\geq 3$) outweigh the steric tissue cost of junction manifolds in the physiological parameter space. The universal persistence of $N=2$ in macroscopic biological reality therefore constitutes a mathematical proof that binary branching cannot be explained by steady-flow metabolic optimisation alone. We identify the missing constraint as the thermodynamic penalty of pulsatile wave-reflection (impedance mismatch) at high-degree nodes: this dynamic mechanism enforces $N=2$ through a network-level variational principle that lies outside the purely static framework developed here, but whose physical necessity is mandated by the present static insufficiency result. We emphasise that this identification is proposed as a resolution of model underdetermination rather than a deductive consequence: the static insufficiency result establishes that the model requires extension, and wave-reflection dynamics constitute the physically motivated candidate for that extension. As an order-of-magnitude estimate, the reactive power fraction in a pulsatile coronary tree scales as $\mathrm{Wo}^{-2}\sim 0.2$ at the aortic root ($\mathrm{Wo}\approx 2.3$), providing a thermodynamic penalty of comparable order to the static cost gap — sufficient to break the degeneracy that the static framework cannot resolve.
Remark 5 (Topological crossover at the capillary limit).
The metabolic order parameter $\kappa\equiv m_{w}/b$ provides a direct physical explanation for the architectural transition from arterial trees to capillary meshes. In macroscopic vessels the active smooth-muscle layer enforces $\kappa\approx 13$. Although this value is insufficient to select $N=2$ under purely static optimisation (Corollary 15), it does generate a thermodynamic preference for small branching integers via the junction penalty $\mathcal{C}_{\mathrm{junctions}}\propto\kappa N^{\alpha_{2}}$ (Theorem 12).
At the capillary level the smooth-muscle layer vanishes entirely, reducing $\kappa\to 1$. The junction penalty—already too weak to enforce $N=2$ in arteries—collapses to the basal endothelial baseline, making higher-order nodes ($N\geq 3$) fully accessible. The emergence of anastomotic capillary meshes is therefore not an anomaly but the exact geometric expression of the system crossing the $\kappa\to 1$ boundary: without the dynamic wave-reflection penalty to enforce $N=2$, and without sufficient junction cost to disfavour $N\geq 3$, reticulated topologies become the natural ground state.
Remark 6 (Interpretation and forward connection).
Proposition 14 provides a structural argument for binary branching rooted in tree architecture. For fixed reach ($M$ terminals), bifurcating trees are the deepest ($G$ maximized at $N=2$), and depth amplifies the cumulative wall investment per unit metabolic cost. The effective stiffness ratio $\kappa_{\mathrm{eff}}$ measures precisely how far the three-term cost function departs from Murray’s homogeneous limit: a larger $\kappa_{\mathrm{eff}}$ generates a steeper non-universality, so $N=2$ is the architecture in which wall-cost selection pressure is strongest.
This structural observation admits a rigorous static completion in the wall-cost-dominated regime. When $\kappa=m_{w}/b\gg 1$, the junction cost overwhelms the tube cost and $\mathcal{C}_{\mathrm{tot}}(N)\approx\mathcal{C}_{\mathrm{junctions}}(N)\propto N^{(1-p)/(5+p)}$, which is strictly increasing in $N$; in this limit, $N=2$ is the unique minimizer of the static cost. For coronary arteries, $\kappa=m_{w}/b\approx 13\gg 1$, so this limit is biologically realized. Together, Theorem 12 (necessary condition: $N<\infty$) and Proposition 14 with the wall-dominated approximation provide a qualitative structural argument for binary branching from wall-cost geometry alone. We note, however, that $\kappa\approx 13$ is pre-asymptotic with respect to the exact threshold $\tau(p)\approx 103$: the wall-dominated limit $\mathcal{C}_{\mathrm{tot}}\approx\mathcal{C}_{\mathrm{junctions}}$ is not yet fully realised at physiological parameters. The asymptotic argument therefore provides a structural insight into the directionality of selection pressure rather than a quantitative demonstration of binary selection—a role filled by Corollary 15. The conclusion is thus robust not through parametric precision but through order-of-magnitude separation: any multiplicative perturbation of $K_{1}$ or $K_{2}$ by a factor less than $10^{2}$ leaves the inequality $K_{2}/K_{1}\ll\tau(p)$ intact.
5.1 Robustness to Non-Isometric Scaling
Theorem 12 assumes the standard self-similar scaling $L\propto r$. In real physiological networks, average arterial length-radius data are well described by a generalized power law $L=k\,r^{\beta}$, where $\beta$ represents the arboreal extension invariant.
Substituting this empirical scaling into the local cost function yields $T(N)=\Phi^{*}(r_{0})\cdot kr_{0}^{\beta}+N\cdot\Phi^{*}(r_{1})\cdot kr_{1}^{\beta}$. In the wall-dominated limit, $\Phi^{*}\propto r^{1+p}$ and $r_{1}=N^{-2/(5+p)}r_{0}$. The combined local length-cost scales as $r^{1+p+\beta}$, leading to a total daughter cost of:
$N\cdot\left(N^{-\frac{2}{5+p}}r_{0}\right)^{1+p+\beta}=N^{1-\frac{2(1+p+\beta)}{5+p}}r_{0}^{1+p+\beta}$ (10)
The strict preference for bifurcations ($N=2$) requires this exponent to be strictly positive, meaning $1>\frac{2(1+p+\beta)}{5+p}$, which simplifies to the condition $\beta<\frac{3-p}{2}$.
For the cardiovascular system where $p\approx 0.77$, the framework strictly preserves $N=2$ for all $\beta<1.115$. This perfectly bounds the physiological range ($\beta\approx 1.0$) observed across species [6], proving that binary branching is structurally robust against macroscopic tissue distortions.
6 Numerical Verification
To ensure no circular reasoning, all parameters are drawn from sources independent of branching-exponent measurements.
Parameters.
[…]
Results.
[…]
Remark 7 (Optimal passive radius and active tone).
Table 2 predicts $r^{*}\in[0.94,1.20]$ mm at $Q=1.3$ mL/s, systematically below the morphometric value $r_{0}=1.5$ mm [6] by 0.30.30.60.6 mm. The present model minimizes the purely passive metabolic cost and therefore predicts the optimal radius absent active tone contributions, not the in-vivo active radius under physiological perfusion pressure.
[…]
Since Theorem 4 and Corollary 6 require only that $B,C>0$, the bounds $(5+p)/2<\alpha^{*}(Q)<3$ are preserved exactly under any positive renormalization of $B$. The predicted passive radius therefore represents a testable lower bound on coronary caliber under maximal pharmacological vasodilation, accessible to standard pressure-myograph or hyperemic flow protocols. This discrepancy is physiologically consistent with active vasomotor tone. While our framework predicts the structural (passive) optimum $r^{*}\approx 1.2$ mm, in vivo coronary arteries exhibit significant basal constriction. A vasomotor contraction of approximately 20% would reconcile our theoretical optimum with the higher morphometric values observed in dilated or fixed specimens.
Table 2: Predicted $\alpha^{*}$ for porcine coronary arteries. No fitting; all parameters from independent literature.
[…]
Experimental [6] $2.7\pm 0.2$
[…]
Figure 1: Scale-dependence of the local branching exponent $\alpha^{*}(Q)$. As flow $Q$ increases, the exponent monotonically approaches the Murray limit ($\alpha=3$) from below. Thicker walls (higher metabolic wall cost $m_{w}$) shift the entire curve toward the wall-dominated lower bound $(5+p)/2\approx 2.885$. The shaded region marks the empirical morphometric range for porcine coronary arteries [6].
Table 3: Cross-network validation. Measured wall scaling exponent $p$ and the associated single-term limit $\alpha=(5+p)/2$. When boundary thickness is governed exactly by vessel radius (e.g. Barlow’s stress formula for internal pressure: $h\propto r$), $p=1$ and the bounds collapse, recovering Murray’s theoretical $\alpha=3$ perfectly.
Network $p$ $\alpha$ Limit $\alpha_{\mathrm{exp}}$
Human pulmonary artery [4] 0.60 2.800 2.7–2.8
[…]
Plant xylem [3] 1.00 3.000 (Murray) $\sim 3$
Engineered pipes (pressure) 1.00 3.000 (Murray) $\sim 3$
The wall-thickness law also allows robust comparisons across highly disparate physical systems (Table 3). Human pulmonary arterial trees exhibit an independent morphometric exponent of $\alpha\approx 2.7$$2.8$ [4]. Their characteristically thinner walls ($p\approx 0.60$) naturally push the theoretical limit lower than the systemic circulation, matching the direction of the empirical shift without requiring parameter fitting. Conversely, for networks designed statically to withstand internal pressure (engineered pipes or certain plant xylem [3]), mechanics dictate $h\propto r^{1}$. Setting $p=1$ forces the bounds to collapse: $(5+1)/2=3\leq\alpha^{*}\leq 3$. The optimization forces Murray’s law to be recovered exactly.
7 Testable Predictions
  1. 1.
    Scale gradient. Since $\lambda\propto r_{0}^{p-1}$ is a strictly decreasing function of $r_{0}$ (and hence of $Q$) for $p<1$, and $\rho^{*}$ is strictly decreasing in $\lambda$ (by implicit differentiation of (6)), $\alpha^{*}(Q)=\ln 2/\ln(1/\rho^{*})$ is strictly increasing in $Q$. Thus $\alpha^{*}(Q)$ monotonically approaches 3 from below as $Q$ increases (Corollary 9). Distal bifurcations (small $Q$) should have slightly smaller $\alpha$ than proximal bifurcations. Predicted range: $\Delta\alpha^{*}<0.015$ across four decades.
  2. 2.
    Wall-cost sensitivity. Thicker-walled vessels (higher $h/r$, i.e. larger $C/B$) should have smaller $\alpha^{*}$, approaching the wall-cost limit $(5+p)/2$.
  3. 3.
    Physical grounding of $\gamma$. Our Theorem 3 gives $\alpha=(4+\gamma)/2$ where $\gamma=1+p$ is the wall-cost exponent. The same formula was obtained by Bennett [1] as $\alpha=(m+4)/2$ with $m$ as a free pricing parameter. Our framework predicts $m=1+p$ for any network whose maintenance cost is dominated by a wall-like sheath. Networks with known $p$ (from histology or materials data) can test this prediction without morphometric fitting. We emphasize that this physical identification $\gamma=1+p$, and hence the Bennett formula $\alpha=(4+\gamma)/2$, holds rigorously only in the two-term limit where either $B=0$ or $C=0$. In the physically complete three-term case ($B,C>0$), Corollary 6 guarantees that no universal exponent exists and the Bennett single-index formula does not apply. This restriction to the two-term class is now fully generalised: Corollary 7 establishes that for any cost function $\Phi=A(Q)r^{-n}+Br^{m}+Cr^{k}$ with $m\neq k$ and $B,C>0$, no universal $\alpha$ exists regardless of the specific transport physics encoded in $A(Q)$.
8 Relation to Prior Work
Murray [9]: two-term cost, $\alpha=3$ (exact). Our Corollary 6 shows this is the only instantiation of the cost-function family (3) with a universal exponent.
Bennett [1]: develops a variational framework for the homogeneous two-term class, deriving generalized Murray scaling, Young–Herring junction geometry, concave (Gilbert-type) flux costs, and a single-index (rigidity index $\chi$) unification showing that the two-term form is the unique quadratic scale-free ledger compatible with simultaneous power-law flux–radius scaling and power-law flux-only concavity [1]. The two-term limit of our Theorem 3 ($\alpha=(4{+}\gamma)/2$ with the exponent $\gamma=1+p$ derived from histological data) coincides with the generalized Murray scaling independently established by Bennett through abstract homogeneity axioms. Our Theorem 4 and Corollary 6 address the regime that lies outside Bennett’s two-term class: incorporating the third biological term ($h(r)\propto r^{p}$, sublinear conduit tissue) breaks Euler homogeneity altogether, producing a rigorously non-universal regime with scale-dependent $\alpha^{*}(Q)$.
Liu & Kassab [8] and Kim & Wagenseil [7] previously incorporated metabolic wall costs into branching models, but explicitly concluded that the standard scaling relations ($\alpha=3$) remained largely invariant. This discrepancy underscores our central thesis: it is specifically the sub-linear allometric scaling of the wall ($p<1$) that breaks the cubic law. Without it, standard three-term energy models natively recover Murray’s limits.
Taylor et al. [16] report optimal Murray exponents of $P=2.15$ (for volumetric flow) and $P=2.38$ (for microvascular resistance) in human epicardial coronary arteries, using a CFD model constrained by pressure-wire measurements. We note that these values lie below our theoretical lower bound $(5+p)/2\approx 2.885$. This is not a contradiction: Taylor et al.’s exponents are hemodynamic optimality exponents—the best-fit $P$ in the relationship $Q\propto D^{P}$ that reproduces in-vivo flow distributions in a clinical population with coronary disease—rather than morphometric diameter-scaling exponents in the sense of Eq. (1). The two quantities coincide only if the tree were exactly Murray-optimal; in diseased or developmentally constrained vascular beds the hemodynamic exponent reflects additional physiological constraints absent from the static cost function. The morphometric exponent most directly comparable to our framework is that of Kassab et al. [6], $\alpha_{\mathrm{exp}}=2.7\pm 0.2$, measured from perfusion-fixed casts of healthy porcine coronary trees, which lies within our predicted range. The sub-$2.885$ values reported by Taylor et al. suggest that pulsatile wave costs introduce an additional architectural constraint not captured by the static cost function analyzed here.
Smink et al. [14]: extend Murray’s two-term optimization to turbulent flows, non-Newtonian fluids, and rough-wall channels, deriving the exponent $x$ in $Q\propto r^{x}$ across the full Reynolds-number parameter space. Their framework covers the fluid-rheology dimension of the optimization problem, while ours covers the biological wall-tissue dimension. Neither framework addresses the signal-transport dimension (such as pulsatile wave reflection in arteries or electrical signal attenuation in neurons), which we propose as the final axis required for a complete universal classification.
West, Brown & Enquist [17]: minimize total network resistance over a self-similar infinite tree, obtaining Kleiber’s $3/4$ law. Our optimization is at the single-vessel level; the network structure enters only through Murray’s branching condition (Lemma 2).
Rall [12]: the branching exponent $\alpha=3/2$ for electrotonic signal propagation in neurons is derived from impedance-matching in cable theory ($Z\propto r^{-3/2}$), not from cost minimization. Corollary 6 therefore does not contradict Rall: the two universality results arise from physically distinct mechanisms (energy minimization vs. impedance matching) and cover non-overlapping parameter ranges. However, this mathematical structure strongly suggests that a generalized Lagrangian unifying metabolic maintenance cost with signal attenuation impedance could natively bridge vascular and neural morphologies.
9 Conclusion
We have shown that the three-term cost function (2) leads to several rigorous results unavailable in the existing two-term Murray framework (generalized by Bennett [1] via a single-index formalism): unique optimal radius (Theorem 1), equivalence of local optimization with global tree law (Lemma 2), exact single-term classification $\alpha=(4+\gamma)/2$ with physical grounding of $\gamma$ (Theorem 3), strict non-universality of the three-term branching exponent extending to asymmetric flows (Theorem 4, Corollary 5), breaking of structural degeneracy to bound the optimal topology to small branching integers, excluding star-like multifurcations (Theorem 12), and tight theoretical bounds on the optimal symmetric bifurcation angle, independently confirmed by three-dimensional morphometric data ($74.9^{\circ}<2\theta^{*}<80.2^{\circ}$), parameterizing the deviation from Murray’s geometry (Theorem 11). The central prediction $\alpha^{*}\in[2.90,2.94]$ from independently measured parameters measurably reduces the gap between Murray’s cubic law and cardiovascular data.
The static attractor as a diagnostic bound.
The purely static wall-tissue mechanism establishes a strict theoretical expectation of $\alpha^{*}\approx 2.90$ for porcine coronary arteries, demonstrating that $\approx 2.90$—not Murray’s 3.03.0—is the structural baseline for purely dissipative biological networks.
This reframing transforms the residual gap with empirical cardiovascular data ($\alpha_{\exp}\approx 2.5$$2.7$) from a modelling discrepancy into a precise physical diagnostic. The macroscopic depression of the empirical exponent below the $2.90$ static attractor serves as an explicit mathematical signature that the cardiovascular tree is subject to physical constraints invisible to a purely dissipative ledger.
Furthermore, Corollary 15 establishes that static optimisation cannot explain why the vascular tree is binary. Both failures—the exponent gap and the topological selection—signal that the final physiological architecture must emerge from a variational principle that balances extensive metabolic dissipation against dimensionless wave-reflection penalties. Formulating such a unified network-level framework represents the necessary definitive extension of the static analysis presented here.
Data and Code Availability.
All computation scripts, numerical data, and figure-generation code are openly and unconditionally available at https://github.com/rikymarche-ctrl/vascular-networks-theory under the CC BY 4.0 Licence. No access request is required.